OpenStudy (anonymous):

I need to find all the real numbers for (2x-1)/(4+x) <= 2

OpenStudy (anonymous):

what I tried was to move the 2 so that I would get (2x-1)/(4+x)-2 <= 0 and from there I got -9/(4+x) <=0

OpenStudy (anonymous):

\[\frac{2x-1}{4+x} \le 2\]Step 1: multiply both sides by 4+x. What do you get?

OpenStudy (anonymous):

yea, I'm pretty sure that I'm not allowed to do that

OpenStudy (anonymous):

Awww... It seems like...... I'm on the wrong track...

jhonyy9 (jhonyy9):

@invidos so because there are x on denominator my opinion that in the first step you need make the restriction for what value not can getting newer x can you do it ?

OpenStudy (anonymous):

x != -4

jhonyy9 (jhonyy9):

yes x not can being never equal -4 so because do you can tel me now why ?

OpenStudy (anonymous):

because you cannot divide by zero and 4+(-4) = 0 or am I missunderstanding you'r question?

jhonyy9 (jhonyy9):

yes do you know it right but with ,,mathanswer" so because than this fraction will be undefined ok ?

OpenStudy (anonymous):

yea

jhonyy9 (jhonyy9):

so now can you continue it ?

OpenStudy (anonymous):

ehm, so I should just multiply both sides and than say that it is undefined when x =-4?

jhonyy9 (jhonyy9):

yes but this x not can being equal -4 this wann being just the first step when you need to solve inequality like in this case ok ? so than now you can continue it how you have wrote little up @RolyPoly

OpenStudy (anonymous):

Nope, the 2x cancels out on both sides!

OpenStudy (anonymous):

x=-4 is a ''critical point'' When x> -4, this inequality is true. When x<-4, this inequality is not true.

OpenStudy (anonymous):

ehm, I am feeling kind of slow right now, but when multiplying both sides with (4+x) you get -1<=8 that might be true but it doesn't give anything

OpenStudy (anonymous):

ok, but isn't there a "smart" way to solve that? I undestand how you got that x>-4 but it seems like you just tested you way there

OpenStudy (anonymous):

Because I don't think you can 'solve' it using 'normal' way...

jhonyy9 (jhonyy9):

so what will get you if you add to both sides (-2) ?

OpenStudy (anonymous):

2x-1-8-2x <= 0 if you then multiply both sides with 4+x

OpenStudy (anonymous):

the 2x still cancels out!

OpenStudy (anonymous):

I guess that you could rewrite everything as (-9)/(4+x) <= 0 which would make it alot easier to se that it cant be smaller than -4

OpenStudy (anonymous):

You guys are smart :D